OpenStudy (anonymous):

How do you write an equation that's in slope intercept or point slope form in standard form( Ax+By=C)?

OpenStudy (midhun.madhu1987):

The slope-intercept form of a lne is given by: y = mx + c where m = slope and c = y-intercept

OpenStudy (midhun.madhu1987):

Point-Slope form of a line is given by: y - y1 = m( x - x1)

OpenStudy (ddcamp):